# From Imitation to Refinement – Residual RL for Precise Assembly

### v3

@@ Abstract
Advances in behavior cloning (BC), like action-chunking and diffusion, have enabled impressive capabilities. Still, imitation alone remains insufficient for learning reliable policies for tasks requiring precise aligning and inserting of objects, like assembly. Our key insight is that chunked BC policies effectively function as trajectory planners, enabling long-horizon tasks. Conversely, as they execute action chunks open-loop, they lack the fine-grained reactivity necessary for reliable execution. Further, we find that the performance of BC policies saturates despite increasing data. Reinforcement learning (RL) is a natural way to overcome BC’s limitations, but it is not straightforward to apply directly to action-chunked models like diffusion policies. We present a simple yet effective method, ResiP (Residual for Precise Manipulation), that sidesteps these challenges by augmenting a frozen, chunked BC model with a fully closed-loop residual policy trained with RL. The residual policy is trained via on-policy RL, addressing distribution shifts and introducing reactive control without altering the BC trajectory planner. Evaluation on high-precision manipulation tasks demonstrates strong performance of ResiP over BC methods and direct RL fine-tuning.
Videos, code, and data are available at https://residual-assembly.github.io.
@@ I Introduction
Robotic manipulation tasks that require both long-horizon planning and high-precision control, such as assembly, remain a significant challenge in robotics [2, 3, 4, 5, 6]. As an example, consider the furniture assembly task depicted in LABEL:fig:overview (Left), where a robot performs assembly by sequentially grasping and then re-orienting a table leg into the correct pose, performing a precise insertion, and finally screwing the leg in place. This representative task involves executing a specific sequence of skills over hundreds of timesteps, with each stage dependent on the successful completion of the previous. Small imprecisions anywhere in the chain can lead to task failure, underscoring the need for reliable execution of diverse skills in the sequence.
@@
Behavior cloning (BC) is a popular approach for teaching robots various manipulation skills [7, 8, 9, 10, 11, 12, 13, 14, 6, 15]. Recent innovations in BC, such as diffusion models [16, 17, 18, 19] and action chunking [20, 6, 16, 18], have improved learning long-horizon, complex behaviors from demonstrations [21, 15]. However, our analysis shows fundamental limitations in BC when applied to tasks requiring high precision. For example, in the insertion task shown in LABEL:fig:overview (Left), a diffusion-based BC model achieves a success rate of $\sim$50with a handful of demonstrations, but performance plateaus even when scaling up to large datasets (e.g., 100,000 demonstrations), reaching only about $\sim$80success, as seen in LABEL:fig:overview (Right). Other recent work finds similar performance saturation [22, 23, 24].
@@
We hypothesize that this saturation occurs because of two issues: First, BC methods suffer from compounding errors due to distributional shifts; the policy operates on states that deviate from those seen during training, leading to errors that accumulate over time [22]. While approaches like Dataset Aggregation (DAgger) [1] can mitigate this error mode by collecting on-policy data, such approaches assume an expert who can be queried on demand, which is usually unavailable. Second, modern BC policies, particularly ones employing action chunking, act more like open-loop “planners” than reactive controllers. Chunked BC policies are limited in compensating for noise and disturbances (i.e., non-reactive) because each chunk runs in an open loop. In tasks such as furniture assembly, specific “bottleneck” states, like insertions, require precise actions at specific time steps. If these critical moments fall within an action chunk, the BC trajectory planner cannot make real-time adjustments to compensate for inaccuracies in the planned actions. This inability to correct actions can lead to task failure in high-precision tasks.
Reinforcement Learning (RL) promises to autonomously learn corrective behaviors through interaction and exploration guided by a reward function [25, 26]. RL has been effective in training robust reactive controllers that can handle a wide array of tasks [27, 28, 29, 30, 31]. However, applying RL to solve complex, long-horizon, and precise tasks like robotic assembly requires extensive reward engineering that can be prohibitively difficult and time-consuming [32]. RL requires vast data, which is costly to collect, particularly in real-world settings.
We present a way to leverage the strengths of BC and RL that reduces human effort in obtaining reliable policies by building on recent advances in BC methods and sim-to-real RL training. BC learns a “planner” that predicts trajectory segments (i.e., an action chunk executed open loop) conditioned on the agent’s observations from a small set of demonstrations, providing a skeleton for RL training. RL augments the BC-learned planner with a closed-loop controller capable of fine-grained corrections that mitigate the limitations of open-loop action chunk execution and collects on-policy data to ensure reliable plan execution. While combining BC and RL has a long history [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], the recent advancements in BC architectures present new challenges for fine-tuning the learned policy with RL. The structure of diffusion models (iterative refinement) and action-chunked policies (large effective action spaces) complicate the application of standard RL algorithms, often leading to instability or requiring significant architectural modifications [46, 15, 47].
We propose a simple yet effective method to combine the strengths of modern chunked BC and on-policy RL controllers, called ResiP (Residual for Precise Manipulation), shown in Fig. 1. Our key insight is recognizing that modern BC policies function as trajectory planners that generate action sequences but crucially lack the fine-grained reactivity and corrections required for reliable execution. We introduce closed-loop corrections by augmenting a frozen, chunked BC model with a small, single-step residual policy [48] trained via on-policy RL. The residual policy observes each state and action and predicts a corrective action to be added to the action predicted by the BC model. This design sidesteps the difficulties of fine-tuning complex BC models with RL while addressing BC’s limitations. First, the residual RL policy operates on-policy, learning reactive behaviors directly through interaction with the environment, thus overcoming issues of distributional shift. Second, providing closed-loop control enables real-time adjustments during task execution, improving robustness and precision, especially in bottleneck stages like insertions. Finally, the residual training is supervised with a sparse task-completion reward, avoiding reward engineering.
We evaluate ResiP on a range of high-precision manipulation tasks from the challenging FurnitureBench [5] and Factory [49] benchmarks and two hand-designed tasks. Our experiments demonstrate that ResiP outperforms pure BC methods and traditional RL fine-tuning, for example, achieving success rates up to $\sim$98starting from 50 demonstrations for the one_leg task. At the same time, increasing data for BC saturates at $\sim$80success with 100K demonstrations for the same task. We also find that ResiP, combined with straightforward applications of teacher-student distillation, visual domain randomization, and co-training, we can transfer policies from simulation to physical robot hardware using RGB observations.
@@
Figure 1: Overview of ResiP. A pretrained chunked base policy predicts an action chunk of $T_{a}$ future actions. For every timestep, the residual model observes the current state $s_{t}$ and predicted base action $a_{t}^{\text{base}}$ and corrects.
II Problem Setup and Approach Overview
@@
Figure 2: Initial and goal states for our 6 tasks. The first three tasks, one_leg, round_table, and lamp, come from the FurnitureBench [5] task suite. mug-rack we created by scanning objects and importing them into the simulation. peg-in-hole is from the Factory [49] task suite. biman-insert we created using the simulation-based teleoperation system DART [50]. Our task suite exhibits diverse challenges like long horizons, tight tolerances, bimanual control, and multi-modal success criteria.
Assumptions and System Components
@@
Our method applies to tasks that can be simulated using already available assets (e.g., CAD models) or using assets that can be obtained via scans of real-world environments [30, 51]. While our method is general, we empirically verify it on a set of challenging assembly tasks of one_leg, round_table, and lamp taken from the FurnitureBench [5] task suite, the peg-in-hole task from [49], a custom mug-hanging task we call mug-rack, and a custom bimanual precise insertion task we call biman-insert. Visualizations of initial state distributions and detailed task descriptions are provided in App. B, and task rollouts can be seen on the project website. Multi-part assembly interactions are simulated using the SDF-based collision geometry representations provided as part of the Factory [49] extension of NVIDIA’s Isaac Gym simulator [52]. The biman-insert task is implemented using the MuJoCo simulator [53], and demos are provided using the DART teleoperation system from [50].
@@
We define a sparse task-completion reward set to 1 for each task when a pair of parts has been fully assembled and 0 otherwise. Many of our tasks are long horizon (up to $\sim$750-1000 steps at 10Hz) and require sequencing of behaviors such as 6-Degree-of-Freedom (DoF) grasping, reorientation, insertion, and screwing (see, e.g., LABEL:fig:overview (Left)).
Preliminaries
Each task is formulated as a discrete-time sequential decision-making problem. At every time-step $t$, the robot receives an observation $o_{t}\in\mathcal{O}$ corresponding to the world and robot state $s_{t}\in\mathcal{S}$ (available only in simulation) and uses it to predict an action $a_{t}$ that is then executed.
The policy operates at 10Hz on a 7-DoF Franka Emika Panda robot arm for all tasks but biman-insert, which operates at 50Hz on two Franka Panda arms. The action space consists of the desired end-effector pose $\mathbf{T}{}^{\text{des}}\in$ $\text{SE}(3)$, which includes both position and orientation and a binary gripper command for opening/closing the parallel-jaw gripper. These desired end-effector poses are converted to joint position targets using differential inverse kinematics [54], then tracked using a low-level PD controller running at 1KHz with manually specified stiffness and dampening parameters.
In the simulated task variants, the system state space $\mathcal{S}$ contains the same end-effector pose $\mathbf{T}$, spatial velocity $\mathbf{V}$, and gripper width $w_{\text{g}}$, along with the 6-DoF poses of all the parts in the environment $\{\mathbf{T}{}^{\text{part}_{i}}\}_{i=1}^{\text{num\_parts}}$. In addition to the state and observation, the simulated agent receives a binary reward signal, indicating whether a pair of assembly parts has achieved their required geometric alignment. For example, in the lamp task, the policy receives two binary rewards: the first when the bulb is fully screwed in and the second when the lamp shade is successfully placed. The real world observation space $\mathcal{O}$ contains the robot end-effector pose $\mathbf{T}{}\in$ $\text{SE}(3)$, robot end-effector spatial velocity $\mathbf{V}\in\mathbb{R}^{6}$, the gripper width $w_{\text{g}}$, and RGB images from a fixed front-view camera ($I^{\text{front}}\in\mathbb{R}^{h\times w\times 3}$) and a wrist-mounted camera ($I^{\text{wrist}}\in\mathbb{R}^{h\times w\times 3}$), each with unknown camera poses.

### v4

@@ Abstract
Recent advances in Behavior Cloning (BC) have made it easy to teach robots new tasks. However, we find that the ease of teaching comes at the cost of unreliable performance that saturates with increasing data for tasks requiring precision. The performance saturation can be attributed to two critical factors: (a) distribution shift resulting from the use of offline data and (b) the lack of closed-loop corrective control caused by action chucking (predicting a set of future actions executed open-loop) critical for BC performance.
Our key insight is that by predicting action chunks, BC policies function more like trajectory “planners” than closed-loop controllers necessary for reliable execution. To address these challenges, we devise a simple yet effective method, ResiP (Residual for Precise Manipulation), that overcomes the reliability problem while retaining BC’s ease of teaching and long-horizon capabilities. ResiP augments a frozen, chunked BC model with a fully closed-loop residual policy trained with reinforcement learning (RL) that addresses distribution shifts and introduces closed-loop corrections over open-loop execution of action chunks predicted by the BC trajectory planner. Videos, code, and data: residual-assembly.github.io.
@@ I Introduction
Many robotic manipulation tasks, such as assembly, require both long-horizon planning and high-precision control, which remains a significant challenge [2, 3, 4, 5, 6]. As an example, consider the furniture assembly task depicted in LABEL:fig:overview (Left), where the robot needs to perform a sequence of steps: grasp, then re-orient the table leg into the correct pose, insert the leg, and finally screw the leg into place. This representative long-horizon task spans several task stages over hundreds of timesteps, each dependent on the successful completion of the previous. During critical moments such as insertions, called “bottleneck states,” even slight imprecisions can compound and result in task failure, underscoring the need for reliable execution of each phase.
@@
Behavior cloning (BC) is a popular approach for teaching robots various manipulation skills [7, 8, 9, 10, 11, 12, 13, 14, 6, 15, 16]. Recent innovations in BC, such as diffusion models [17, 18, 19, 20] and action chunking (predicting a sequence of future actions) [21, 6, 17, 19], have enabled learning long-horizon, complex behaviors from demonstrations [22, 15]. However, our analysis shows fundamental limitations in BC when applied to tasks requiring high precision: performance saturates with increasing data. For example, the success of a diffusion-based BC model on an insertion task shown in LABEL:fig:overview (Left) plateaus at $\sim$80even with 100,000 demonstrations (LABEL:fig:overview; right). Recent work finds similar performance saturation [23, 24, 25].
@@
We hypothesize that the performance saturation results from two issues: (i) compounding errors originating from distribution shift as the policy operates on states that increasingly deviate from those seen during training [23]. (ii) The use of action-chunking which enables long-horizon control at the cost of reactivity necessary for reliable execution by rolling out each action chunk open-loop [26]. Therefore, we posit that chunked BC policies are best considered “planners” rather than reactive controllers. For instance, in the one_leg assembly task shown in LABEL:fig:overview (left), specific bottleneck states, like insertions, require precise actions at specific time steps. If these critical moments fall within an action chunk, the BC policy cannot make real-time adjustments to compensate for disturbances during execution or inaccuracies in the planned actions.
Reinforcement Learning (RL) is a standard solution to the suboptimal performance of BC polices [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. RL fine-tuning effectively overcomes the problem of distribution shifts by generating training data for states visited by the policy. However, recent advancements in BC architectures present new challenges for direct RL fine-tuning. The main issues are that the structure of diffusion models (iterative refinement) and action-chunked policies (resulting in a large action space) make standard RL algorithms unstable (see Sec. IV-A) or require significant architectural modifications [40, 15, 41].
Instead of invoking RL, one can overcome distribution shift by leveraging an expert and training with supervised learning such as the Dataset Aggregation (DAgger) [1] algorithm. Our experiments show with roughly  2000 demonstrations, a DAgger style approach achieves a 90% success rate on the one_leg task (LABEL:fig:overview)—a significant improvement over pure BC but still 8% below the expert. One challenge in using DAgger is that it assumes an expert who can be queried on demand, which is usually unavailable. However, another significant challenge is the performance saturation of DAgger, which we attribute to the lack of reactivity because each action chunk is executed in an open-loop fashion.
Our key insight is that action-chunked BC policies function more like “trajectory planners” than reactive controllers. To mitigate both distribution shift and lack of closed-loop control, we use a simple method of augmenting a frozen, chunked BC model with a small, single-step residual policy [42, 43, 44, 45, 46, 47, 48, 49] trained via on-policy RL (Fig. 1), termed ResiP (Residual for Precise Manipulation). Unlike prior residual frameworks, the BC policy generates trajectory segments at coarse temporal resolution ($\sim$1 second), while the residual predicts closed-loop corrections at each control step ($\sim$0.1 second).
ResiP addresses both challenges faced by BC methods: the RL-trained residual overcomes distribution shifts and by closing the sensorimotor loop more frequently it lends reactivity by enabling precise adjustments that chunked policies cannot provide. Using only a sparse task-completion reward, this approach achieves a 98% success rate starting with just 50 expert demonstrations on the one_leg task—an 18 percentage point improvement over BC with 100K demonstrations.
Since our method relies on training RL policies, our learning pipeline heavily relies on simulation. Applying our pipeline requires constructing a digital twin of the real-world use case (i.e., real-to-sim [16]), training with ResiP, and finally transferring the trained policy to the real world. While our focus in this paper is on ResiP and highlighting the performance saturation of purely BC methods, our experiments demonstrate the entire pipeline on several challenging assembly tasks.
@@
Figure 1: Overview of ResiP. A pretrained chunked base policy predicts an action chunk of $T_{a}$ future actions. For every timestep, the residual model observes the current state $s_{t}$ and the predicted base action $a_{t}^{\text{base}}$ and predicts correction.
II Method
Our method applies to any task that can be simulated using already available assets (e.g., CAD models) or by scanning real-world objects [16, 50]. While this paper focuses on assembly tasks for which CAD models are already available for constructing simulation environments, our method applies to many tasks of practical interest. Note that our method requires CAD models only for training and not during deployment.
Our approach from task definition to deployable vision policy consists of three key components (see Fig. 2 for an overview). First, we train a base policy using behavior cloning (BC) on a small set of demonstrations (Sec. II-B) in simulation. Then, we improve this policy’s precision by training a residual component with reinforcement learning that makes closed-loop corrections to the base policy’s actions (Sec. II-C). Finally, we learn a real-world deployable policy with policy distillation and co-training with a few real-world demonstrations (Sec. II-D).
II-A Problem Formulation
We formulate the target task as a discrete-time sequential decision-making problem. At each timestep $t$, the robot receives an observation $o_{t}\in\mathcal{O}$ corresponding to the underlying state $s_{t}\in\mathcal{S}$ of the robot and environment. Given the observation, the policy outputs an action $a_{t}\in\mathcal{A}$ that is executed in the environment. The action space $\mathcal{A}$ consists of end-effector poses in SE(3) and gripper commands. The state space includes the robot’s configuration (end-effector pose, velocity, and gripper state) and the poses of all objects in the scene. Task completion is defined through sparse task completion rewards. We define this reward as an instantaneous signal that provides 1 at the timestep when a pair of objects achieves a pre-defined relative pose corresponding to successful assembly and 0 otherwise, including after the alignment is maintained. See more details about the environments in App. B-A.
II-B Base Policy Learning via Behavior Cloning
[…]
II-C Reactive Control via ResiP
Given the initial chunked base policy $\pi_{\text{base}}$ obtained by BC described above, we want to improve the policy to overcome the issues of distribution shift and the lack of reactivity. One way to mitigate the adverse effects of distribution shifts is to fine-tune the BC-trained policy with RL [7, 8, 9, 10, 11, 12, 13, 14, 6, 15, 16]. We side-step the complications of direct RL fine-tuning of action-chunked policies (see discussion in Sec. B-A) by training a residual [42, 43, 44, 45, 46, 47, 48, 49] Gaussian Multi-Layer Perceptron (MLP) [51] policy $\pi_{\text{res}}$ using PPO [52]. Our key design choice is that while this residual framework can address distribution shift through either chunk-level or timestep-level corrections (see Sec. III-D4 for analysis), the latter enables reactivity helpful for precise manipulation.
For each timestep $i=0,...,T_{a}-1$ within the base policy’s action chunk, we form the residual’s observation by concatenating the full simulation state (robot proprioceptive information and object poses) with the base policy’s predicted action, $s_{t+i}^{\text{res}}=[s_{t+i},a^{\text{base}}{t+i}]$. The residual policy then produces a corrective action $a_{t+i}^{\text{res}}\sim\pi_{\text{res}}(\cdot|s_{t+i}^{\text{res}})$ that modifies the base action: $a_{t+i}=a_{t+i}^{\text{base}}+a_{t+i}^{\text{res}}$. The resulting fine-tuned policy is a combination of the pre-trained BC policy $\pi_{\text{base}}$ and the correction policy $\pi_{\text{res}}$, denoted $\boldsymbol{\pi}$. As demonstrated in Sec. IV-B3, this per-timestep correction capability is crucial for achieving reliable performance in precision-critical tasks.
During training, as in any RL fine-tuning, one needs to adjust the exploration noise to provide sufficient stochasticity to discover new behaviors while maintaining enough precision for task success (see App. J-C for detailed analyses of design choices). The residual policy uses orthogonal initialization [53] with a small gain factor on the final layer, ensuring uncorrelated initial corrections centered around zero—a design choice that aligns with our assumption that the base policy’s errors are unbiased in expectation. Since most task failures stem from slight imprecisions in the base policy’s actions, this architecture naturally leads to learning small, targeted corrections around the base policy’s behavior.
Figure 3: Photorealistic rendering and domain randomization for sim-to-real transfer. In the first row, we show our nominal experiment setting of rendering parts in the original white color while varying the position, brightness, and hue of the lighting in the scene. In the bottom row, we also introduce variations in the part colors to see how that can help the final policy adapt to unseen part appearances without collecting any additional data.
The policy observation for simulation training, $\mathcal{S}$ contains the 6 DoF end-effector pose $\mathbf{T}$, spatial velocity $\mathbf{V}$, and gripper width $w_{\text{g}}$, along with the 6-DoF poses of all the parts in the environment $\{\mathbf{T}{}^{\text{part}_{i}}\}_{i=1}^{\text{num\_parts}}$.
@@ II-C Reactive Control via ResiP
Figure 4: Initial and goal states for our 6 tasks. The first three tasks, one_leg, round_table, and lamp, are from the FurnitureBench [5] task suite. mug-rack we created by scanning objects and importing them into the simulation. peg-in-hole is from the Factory [54] task suite. biman-insert we created using the simulation-based teleoperation system DART [55]. Our task suite exhibits diverse challenges like long horizons, tight tolerances, bimanual control, and multi-modal success criteria.
II-D Sim-to-Real Transfer
We treat $\pi_{res}$ trained in simulation using privileged state information, $\mathcal{S}$, as a teacher policy $\pi_{\text{teacher}}$ to distill into a student policy $\pi_{\text{student}}$ that takes as input sensory observations and can therefore be deployed in the real world [56, 57, 58, 16]. The real world observation space $\mathcal{O}$ contains the robot end-effector pose $\mathbf{T}{}\in$ $\text{SE}(3)$, robot end-effector spatial velocity $\mathbf{V}\in\mathbb{R}^{6}$, the gripper width $w_{\text{g}}$, and RGB images from a fixed front-view camera ($I^{\text{front}}\in\mathbb{R}^{h\times w\times 3}$) and a wrist-mounted camera ($I^{\text{wrist}}\in\mathbb{R}^{h\times w\times 3}$), each with uncalibrated camera poses.
For teacher-student distillation, we collect a dataset of successful trajectories $\mathcal{D}_{\text{synth}}=\{\tau_{\text{synth},1},...,\tau_{\text{synth},N}\}$, $\tau_{\text{synth},i}=\{(s_{1},a_{1}),...,(s_{T},a_{T})\}$, from $\pi_{teacher}$ initialized with a diverse set of initial object poses. To bridge the sim-to-real gap, we convert the trajectory dataset with only state information ($\mathcal{S}$) into trajectories with sensory observations ($\mathcal{O}$) by re-rendering the trajectories into realistic-looking image observations (see Fig. 3 for examples). This rendering process allows us to introduce visual variations—including object colors, textures, lighting conditions, and camera perspectives—that are not easily obtainable with standard image augmentations or real-world data collection (see App. C for examples and details). We denote this rendered dataset as $\mathcal{D}_{\text{synth-render}}$.
$\pi_{\text{student}}$ is represented with a Diffusion Policy architecture that uses ResNet18 [59] vision encoder pre-trained on robotic manipulation data [60] (see App. A-A3 for implementation details). To minimize the sim-to-real gap, we combine a small set of real-world demonstrations $\mathcal{D}_{\text{real}}$ (10-40 trajectories) collected directly on the robot with synthetic rendered trajectories from $\mathcal{D}_{\text{synth-render}}$ (approximately 400 demonstrations) [16]. This synthetic-to-real data ratio provides increased domain coverage while not washing out real-world demonstrations. The real demonstrations contain only RGB observations and no ground-truth pose information. This combined dataset $\mathcal{D}_{\text{real}}\cup\mathcal{D}_{\text{synth-render}}$ is used to train the student policy $\pi_{\text{student}}$ with the same BC loss as in the simulation-based experiments.
III Experimental Setup
III-A Tasks and Environment
@@
We evaluate our method on a set of challenging assembly tasks of one_leg, round_table, and lamp from the FurnitureBench [5] task suite, the peg-in-hole task from [54], a custom mug-hanging task we call mug-rack, and a custom bimanual precise insertion task we call biman-insert. Examples of all tasks are shown in Fig. 4. Visualizations of initial state distributions and detailed task descriptions are provided in App. B, and task rollouts can be seen on the website.
Multi-part tight-tolerance assembly interactions are simulated using the SDF-based collision geometry representations provided as part of the Factory [54] extension of Nvidia’s IsaacGym simulator [61], with demos collected using a SpaceMouse. The biman-insert task is implemented using the MuJoCo simulator [62], with the demos provided using the DART teleoperation system from [55].
@@
We define a simple, sparse task-completion reward set to 1 for each task when a pair of parts has been fully assembled and 0 otherwise. For example, in the lamp task, the policy receives two binary rewards: the first when the bulb is fully screwed in and the second when the lamp shade is successfully placed. Some of our tasks are long horizon (up to $\sim$750-1000 steps at 10Hz) and require sequencing of behaviors such as 6-Degree-of-Freedom (DoF) grasping, reorientation, insertion, and screwing (see LABEL:fig:overview; Left).